# Sup-norm bounds for Jacobi cusp forms

### v2

@@ 4. Sup-norm bounds for Jacobi cusp forms
For $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, let $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ be a Jacobi cusp form of weight $k$ and index $m$ for the full modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, which is normalized with respect to the Petersson inner product defined by (12). We now aim at bounding the quantity
$\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}.$
For this we will employ the decomposition (13), namely
$f(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z),$
where the functions $\varphi_{\mu}$ are cusp forms of weight $(k-\frac{1}{2})$ with respect to the finite index subgroup $\Gamma_{1}=\Gamma_{0}(4m)$ of $\Gamma_{0}$ and the theta functions $\vartheta_{\mu,m}$ are defined in (14).
For a fixed $\tau=\xi+i\eta\in\mathbb{H}$, consider the elliptic curve $E_{\tau}\coloneqq\mathbb{C}/\Lambda_{\tau}$ with $\Lambda_{\tau}\coloneqq\mathbb{Z}\oplus\,\tau\mathbb{Z}$. Let $O_{\tau}$ denote the identity element of $E_{\tau}$, when considered as an abelian group with $\oplus_{\tau}$ denoting the group operation, and let
$[2]\colon E_{\tau}\longrightarrow E_{\tau}$
be multiplication by $2$, given by the assignment $z\mapsto 2z\coloneqq z\oplus_{\tau}z$, which is an isogeny of degree $4$. Let $\mathcal{M}_{\tau}$ be the line bundle associated to the divisor $O_{\tau}$. Then, the theorem of the cube gives the isomorphism
$[2]^{\ast}\mathcal{M}_{\tau}\cong\mathcal{M}_{\tau}^{\otimes{4}}.$
We then put $\mathcal{L}_{\tau}\coloneqq\mathcal{M}_{\tau}^{\otimes{2}}$ and find that the theta functions $\vartheta_{\mu,m}(\tau,\cdot)$ ($\mu=0,\ldots,2m-1$) arise as global holomorphic sections of the line bundle $\mathcal{L}_{\tau}^{\otimes{m}}$.
The pointwise norm of $\vartheta_{\mu,m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ at the point $z=x+iy\in E_{\tau}$ (identifying $E_{\tau}$ with its universal cover $\mathbb{C}$) is given by the following formula
(46) $\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\coloneqq|\vartheta_{\mu,m}(\tau,z)|^{2}\,\sqrt{\eta}\,e^{-\frac{4\pi my^{2}}{\eta}}.$
Let $\mu_{\mathrm{eucl}}$ denote the Euclidean metric on $E_{\tau}$; at the point $z=x+iy\in E_{\tau}$, it is given by the formula
(47) $\mu_{\mathrm{eucl}}(z)=\frac{i}{2}\cdot\frac{\mathrm{d}z\wedge\mathrm{d}\overline{z}}{\eta}=\frac{\mathrm{d}x\wedge\mathrm{d}y}{\eta}.$
Furthermore, the $\mathrm{L}^{2}$-inner product of $\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu^{\prime},m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ is given by the formula
$\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu^{\prime},m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\mathcal{L}_{\tau}^{\otimes{m}}}$ $\coloneqq\int\limits_{E_{\tau}}\vartheta_{\mu,m}(\tau,z)\,\overline{\vartheta_{\mu^{\prime},m}(\tau,z)}\,\sqrt{\eta}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\mu_{\mathrm{eucl}}(z)$
(48) $=\int\limits_{0}^{\eta}\int\limits_{0}^{1}\vartheta_{\mu,m}(\tau,z)\,\overline{\vartheta_{\mu^{\prime},m}(\tau,z)}\,\sqrt{\eta}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}x\wedge\mathrm{d}y}{\eta}.$
The $\mathrm{L}^{2}$-norm of $\vartheta_{\mu,m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ induced by the $\mathrm{L}^{2}$-inner product defined above is thus given by
(49) $\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu,m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\mathcal{L}_{\tau}^{\otimes{m}}}.$
We next recall the well-known fact about the orthogonality of the Jacobi theta functions (which was already implicitly used in formula (15)).
@@
The set of Jacobi theta functions $\{\vartheta_{\mu,m}(\tau,\cdot)\}_{\mu=0}^{2m-1}$ constitutes an orthogonal basis for the space $H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ with respect to the $\mathrm{L}^{2}$-inner product (48).
@@
As above, we write $\tau=\xi+i\eta\in\mathbb{H}$ and for $\mu,\mu^{\prime}=0,\ldots,2m-1$ and $n\in\mathbb{Z}$, we define
(50) $a(\mu,n)\coloneqq n-\frac{\mu}{2m}\quad\text{and}\quad b(\mu,n)\coloneqq 2mn-\mu.$
Recalling the definition of the Jacobi theta function (14), we compute for the $\mathrm{L}^{2}$-inner product (48)
$\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu^{\prime},m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}$
(51) $\qquad=\int\limits_{0}^{\eta}\sum_{n\in\mathbb{Z}}\,\sum_{n^{\prime}\in\mathbb{Z}}e^{2\pi im\tau a(\mu,n)^{2}}\overline{e^{2\pi im\tau a(\mu^{\prime},n^{\prime})^{2}}}\bigg{(}\int\limits_{0}^{1}e^{2\pi izb(\mu,n)}\overline{e^{2\pi izb(\mu^{\prime},n^{\prime})}}\,\mathrm{d}x\bigg{)}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}.$
Observing that the inequalities $-2m<\mu-\mu^{\prime}<2m$ imply that the equality
$b(\mu,n)=b(\mu^{\prime},n^{\prime})\quad\Longleftrightarrow\quad 2m(n-n^{\prime})=\mu-\mu^{\prime}$
can only hold if $n=n^{\prime}$ and $\mu=\mu^{\prime}$, we deduce
$\int\limits_{0}^{1}e^{2\pi izb(\mu,n)}\overline{e^{2\pi izb(\mu^{\prime},n^{\prime})}}\,\mathrm{d}x$ $=e^{-2\pi y(b(\mu,n)+b(\mu^{\prime},n^{\prime}))}\int\limits_{0}^{1}e^{2\pi ix(b(\mu,n)-b(\mu^{\prime},n^{\prime}))}\,\mathrm{d}x$
(52) =$=\begin{cases}e^{-4\pi yb(\mu,n)},&\text{if }\mu=\mu^{\prime}\text{ and }n=n^{\prime},\\ 0,&\text{else}.\end{cases}$
This proves the pairwise orthogonality of the $2m$ Jacobi theta functions $\{\vartheta_{\mu,m}(\tau,\cdot)\}_{\mu=0}^{2m-1}$ and thus their linear independence. Now, since we have $\deg(\mathcal{L}_{\tau}^{\otimes{m}})=2m$, the Riemann–Roch Theorem on the elliptic curve $E_{\tau}$ shows that
$\dim H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})=2m,$
from which the claim follows. ∎
Proposition 4.2.
With notations as above, we have for $\mu=0,\ldots,2m-1$, the bound
$\sup_{\tau\in\mathbb{H}}\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=O\bigg{(}\frac{1}{\sqrt{m}}\bigg{)}.$
@@
Let $\tau=\xi+i\eta\in\mathbb{H}$, $z=x+iy\in\mathbb{C}$, and $\mu=0,\ldots,2m-1$. For $\tau\in\mathbb{H}$, we need to bound the $\mathrm{L}^{2}$-norm (49) of $\vartheta_{\mu,m}(\tau,\cdot)\in H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$, namely the quantity
$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=\langle\vartheta_{\mu,m}(\tau,\cdot),\vartheta_{\mu,m}(\tau,\cdot)\rangle_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}.$
Recalling the notation (50) and introducing $a(n)\coloneqq a(\mu,n)$ as well as $b(n)\coloneqq b(\mu,n)$ as a short-hand, we arrive as in (51) at the equality
$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}$
(53) $\qquad=\int\limits_{0}^{\eta}\sum_{n\in\mathbb{Z}}\,\sum_{n^{\prime}\in\mathbb{Z}}e^{2\pi im\tau a(n)^{2}}\overline{e^{2\pi im\tau a(n^{\prime})^{2}}}\bigg{(}\int\limits_{0}^{1}e^{2\pi izb(n)}\overline{e^{2\pi izb(n^{\prime})}}\,\mathrm{d}x\bigg{)}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}.$
Since (by (52))
$\int\limits_{0}^{1}e^{2\pi izb(n)}\overline{e^{2\pi izb(n^{\prime})}}\,\mathrm{d}x=\begin{cases}e^{-4\pi yb(n)},&\text{ if }n=n^{\prime},\\ 0,&\text{ else},\end{cases}$
we arrive at
(54) $\sum_{n\in\mathbb{Z}}\,\sum_{n^{\prime}\in\mathbb{Z}}e^{2\pi im\tau a(n)^{2}}\overline{e^{2\pi im\tau a(n^{\prime})^{2}}}\int\limits_{0}^{1}e^{2\pi izb(n)}\overline{e^{2\pi izb(n^{\prime})}}\,\mathrm{d}x=\sum_{n\in\mathbb{Z}}e^{-4\pi m\eta\big{(}n-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2mn-\mu)}.$
Substituting (54) into (53) and using an integral test, we find the bound
$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}$ $=\sum_{n\in\mathbb{Z}}\int\limits_{0}^{\eta}e^{-4\pi m\eta\big{(}n-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2mn-\mu)}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}$
(55) $\leq\int\limits_{0}^{\eta}\int\limits_{-\infty}^{\infty}e^{-4\pi m\eta\big{(}\nu-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2m\nu-\mu)-\frac{4\pi my^{2}}{\eta}}\,\mathrm{d}\nu\,\frac{\mathrm{d}y}{\sqrt{\eta}}$
(56) $\quad+\int\limits_{0}^{\eta}e^{-\frac{\pi\eta\mu^{2}}{m}+4\pi y\mu-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}.$
Now, we rewrite the exponent of the integrand in (55) in the form
$4\pi m\eta\bigg{(}\nu-\frac{\mu}{2m}\bigg{)}^{2}+4\pi y(2m\nu-\mu)+\frac{4\pi my^{2}}{\eta}$
$\qquad=\bigg{(}\sqrt{\frac{\pi\eta}{m}}(2m\nu-\mu)\bigg{)}^{2}+4\pi y(2m\nu-\mu)+\bigg{(}2\sqrt{\frac{\pi m}{\eta}}y\bigg{)}^{2}$
$\qquad=\bigg{(}\sqrt{\frac{\pi\eta}{m}}(2m\nu-\mu)+2\sqrt{\frac{\pi m}{\eta}}y\bigg{)}^{2}.$
Substituting
$\rho\coloneqq\sqrt{\frac{\pi\eta}{m}}(2m\nu-\mu)+2\sqrt{\frac{\pi m}{\eta}}y$
into (55), we obtain for the inner integral
$\int\limits_{-\infty}^{\infty}e^{-4\pi m\eta\big{(}\nu-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2m\nu-\mu)-\frac{4\pi my^{2}}{\eta}}\,\mathrm{d}\nu=\frac{1}{2\sqrt{\pi m\eta}}\int\limits_{-\infty}^{\infty}e^{-\rho^{2}}\,\mathrm{d}\rho=\frac{1}{2\sqrt{m}}\frac{1}{\sqrt{\eta}}.$
From this, we compute the double integral (55) as
(57) $\int\limits_{0}^{\eta}\int\limits_{-\infty}^{\infty}e^{-4\pi m\eta\big{(}\nu-\frac{\mu}{2m}\big{)}^{2}-4\pi y(2m\nu-\mu)-\frac{4\pi my^{2}}{\eta}}\,\mathrm{d}\nu\,\frac{\mathrm{d}y}{\sqrt{\eta}}=\frac{1}{2\sqrt{m}}\frac{1}{\sqrt{\eta}}\int\limits_{0}^{\eta}\frac{\mathrm{d}y}{\sqrt{\eta}}=\frac{1}{2\sqrt{m}}.$
For the integral (56), we find in a similar way
(58) $\int\limits_{0}^{\eta}e^{-\frac{\pi\eta\mu^{2}}{m}+4\pi y\mu-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}\leq\int\limits_{-\infty}^{\infty}e^{-\big{(}-\sqrt{\frac{\pi\eta}{m}}\mu+2\sqrt{\frac{\pi m}{\eta}}y\big{)}^{2}}\,\frac{\mathrm{d}y}{\sqrt{\eta}}=\frac{1}{2\sqrt{m}}.$
Adding up the bounds (57) and (58), yields
$\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\,\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\leq\frac{1}{\sqrt{m}},$
which proves the claim. ∎
Proposition 4.3.
With notations as above, we have the bound
(59) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=O(\sqrt{m}).$
@@
We start by considering the Bergman kernel $B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)$ associated to the line bundle $\mathcal{L}_{\tau}^{\otimes{m}}$ on the elliptic curve $E_{\tau}$ equipped with the Kähler form $\omega=\mu_{\mathrm{eucl}}$ (see formula (47)). From the bound (17), we recall
$B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)=\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)\big{)}\,m+O(1),$
where $\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)$ is the curvature form of $\mathcal{L}_{\tau}$ at the point $z\in E_{\tau}$. Using equations (16) and (46), the curvature form $\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)$ is given by the formula (identifying again $E_{\tau}$ with its universal cover $\mathbb{C}$ and writing $z=x+iy$)
$\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)$ $=-\frac{i}{2\pi}\partial_{z}\partial_{\overline{z}}\log\|\vartheta_{\mu,1}(\tau,z)\|_{\mathcal{L}_{\tau}}^{2}$
$=-\frac{i}{2\pi}\partial_{z}\partial_{\overline{z}}\bigg{(}-\frac{4\pi y^{2}}{\eta}\bigg{)}=2\mu_{\mathrm{eucl}}(z)=2\omega(z).$
This yields $\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L}_{\tau},\|\,\cdot\,\|_{\mathcal{L}_{\tau}})(z)\big{)}=2$, and therefore we get
(60) $B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)=2m+O(1).$
Setting
$\Theta_{\mu,m}(\tau,z)\coloneqq\frac{\vartheta_{\mu,m}(\tau,z)}{\|\vartheta_{\mu,m}(\tau,\cdot)\|_{\mathrm{L}^{2},\mathcal{L}^{\otimes{m}}}}\qquad(\mu=0,\ldots,2m-1)$
and using Lemma 4.1, we deduce that the set $\{\Theta_{\mu,m}(\tau,\cdot)\}_{\mu=0}^{2m-1}$ constitutes an orthonormal basis for the space $H^{0}(E_{\tau},\mathcal{L}_{\tau}^{\otimes{m}})$ with respect to the $\mathrm{L}^{2}$-inner product (48). By the very definition of the Bergman kernel we thus have
$B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z)=\sum_{\mu=0}^{2m-1}\|\Theta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}$
for any $z\in E_{\tau}$. Finally, using Proposition 4.2 and observing that
$\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}\leq\frac{1}{\sqrt{m}}\sum_{\mu=0}^{2m-1}\|\Theta_{\mu,m}(\tau,z)\|_{\mathcal{L}_{\tau}^{\otimes{m}}}^{2}=\frac{1}{\sqrt{m}}B_{\mathcal{L}_{\tau}^{\otimes{m}}}(z),$
the claim follows by means of (60). ∎

### v3

@@ 4. Sup-norm bounds for Jacobi cusp forms
For $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, let $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ be an $L^{2}$-normalized Jacobi cusp form of weight $k$ and index $m$ for the full modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$. We next aim at bounding the quantity
$\|\phi\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|_{\mathrm{Pet}}.$
By the $\Gamma_{0}\ltimes\mathbb{Z}^{2}$-invariance of the function $\|\phi(\tau,z)\|_{\mathrm{Pet}}$ on $\mathbb{H}\times\mathbb{C}$, it suffices to find a bound for the quantity $\|\phi(\tau,z)\|_{\mathrm{Pet}}$, when $\tau$ is ranging through a fundamental domain $\mathcal{F}_{\Gamma_{0}}$ for $\Gamma_{0}$ and $z$ is ranging through a fundamental domain of the elliptic curve $\mathbb{C}/(\mathbb{Z}\oplus\tau\mathbb{Z})$. Without loss of generality, we can assume in the sequel that
$\mathcal{F}_{\Gamma_{0}}=\{\tau=\xi+i\eta\in\mathbb{H}\,|\,-1/2\leq\xi\leq 1/2,\,|\tau|\geq 1\},$
and, given $\tau=\xi+i\eta\in\mathcal{F}_{\Gamma_{0}}$, that
[…]
Furthermore, in order to arrive at an $L^{\infty}$-bound of the quantity $\|\phi(\tau,z)\|_{\mathrm{Pet}}$, we will make use of the decomposition (13), namely
$\phi(\tau,z)=\sum_{\mu=0}^{2m-1}h_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z),$
where the functions $h_{\mu}$ are cusp forms of weight $(k-\frac{1}{2})$ with respect to the finite index subgroup $\Gamma_{1}=\Gamma_{0,1}(4m)$ of $\Gamma_{0}$ and the theta functions $\vartheta_{\mu,m}$ are defined in (14).
@@
With notations as above, we have the bound
(44) $\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|^{2}_{\mathrm{Pet}}\leq 2m\,\mathrm{Im}(\tau)^{\frac{1}{2}}+O\big{(}m^{\frac{1}{2}}\big{)}$
for $(\tau,z)\in\mathbb{H}\times\mathbb{C}$.
@@
By definition, we have for $(\tau,z)\in\mathbb{H}\times\mathbb{C}$ the equalities
$\vartheta_{\mu,m}(\tau,z)$ $=\sum_{n\in\mathbb{Z}}e^{2\pi im\tau(n-\frac{\mu}{2m})^{2}+2\pi iz(2mn-\mu)}$
$=\sum_{n\in\mathbb{Z}}e^{2\pi imn^{2}\tau-2\pi i\cdot 2mn\cdot\frac{\mu\tau}{2m}+\pi i\frac{\mu^{2}\tau}{2m}+2\pi i\cdot 2mn\cdot z-2\pi i\mu z}$
$=e^{\pi i\frac{\mu^{2}\tau}{2m}-2\pi i\mu z}\,\sum_{n\in\mathbb{Z}}e^{2\pi imn^{2}\tau+2\pi i(z-\frac{\mu\tau}{2m})2mn}$
$=\vartheta_{0,m}\Big{(}\tau,z-\frac{\mu\tau}{2m}\Big{)}\,e^{\pi i\frac{\mu^{2}\tau}{2m}-2\pi i\mu z}.$
Taking absolute values, we get
$|\vartheta_{\mu,m}(\tau,z)|=\Big{|}\vartheta_{0,m}\Big{(}\tau,z-\frac{\mu\tau}{2m}\Big{)}\Big{|}\,e^{-\pi\frac{\mu^{2}\eta}{2m}+2\pi\mu y}.$
This implies the relation
$\|\vartheta_{\mu,m}(\tau,z)\|_{\mathrm{Pet}}$ $=|\vartheta_{\mu,m}(\tau,z)|\,\eta^{\frac{1}{4}}\,e^{-\frac{2\pi my^{2}}{\eta}}$
$=\Big{|}\vartheta_{0,m}\Big{(}\tau,z-\frac{\mu\tau}{2m}\Big{)}\Big{|}\,\eta^{\frac{1}{4}}\,e^{-\pi\frac{\mu^{2}\eta}{2m}+2\pi\mu y-2\pi m\frac{y^{2}}{\eta}}$
$=\Big{|}\vartheta_{0,m}\Big{(}\tau,z-\frac{\mu\tau}{2m}\Big{)}\Big{|}\,\eta^{\frac{1}{4}}\,e^{-\frac{2\pi m}{\eta}(y-\frac{\mu\eta}{2m})^{2}}.$
From this, we derive for $(\tau,z)\in\mathbb{H}\times\mathbb{C}$ the bound
$\|\vartheta_{\mu,m}(\tau,z)\|_{\mathrm{Pet}}$ $=\Big{|}\vartheta_{0,m}\Big{(}\tau,z-\frac{\mu\tau}{2m}\Big{)}\Big{|}\,\eta^{\frac{1}{4}}\,e^{-\frac{2\pi m}{\eta}(y-\frac{\mu\eta}{2m})^{2}}$
$\leq\sum_{n\in\mathbb{Z}}e^{-2\pi mn^{2}\eta-4\pi mn(y-\frac{\mu\eta}{2m})-\frac{2\pi m}{\eta}(y-\frac{\mu\eta}{2m})^{2}}\,\eta^{\frac{1}{4}}$
$=\sum_{n\in\mathbb{Z}}e^{-\frac{2\pi m}{\eta}(n\eta+y-\frac{\mu\eta}{2m})^{2}}\,\eta^{\frac{1}{4}}.$
By an integral test, we arrive for $(\tau,z)\in\mathbb{H}\times\mathbb{C}$ at the bound
$\|\vartheta_{\mu,m}(\tau,z)\|_{\mathrm{Pet}}$ $\leq\bigg{(}1+\int_{-\infty}^{\infty}e^{-\frac{2\pi m}{\eta}(n\eta+y-\frac{\mu\eta}{2m})^{2}}\,\mathrm{d}n\bigg{)}\eta^{\frac{1}{4}}$
$=\bigg{(}1+\frac{1}{\sqrt{2m\eta}}\bigg{)}\eta^{\frac{1}{4}},$
which leads to
$\sum_{\mu=0}^{2m-1}\|\vartheta_{\mu,m}(\tau,z)\|^{2}_{\mathrm{Pet}}\leq 2m\eta^{\frac{1}{2}}+O\big{(}m^{\frac{1}{2}}\big{)}.$
This completes the proof of the lemma. ∎
Lemma 4.2.
With notations as above, let $k\in\mathbb{Z}_{\geq 5}$, let $\Gamma\subseteq\Gamma_{0}$ be a subgroup of finite index, whose stabilizer subgroup $\Gamma_{i\infty}$ is generated by $\big{(}\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\big{)}$, and let $f\in S_{k,\chi}(\Gamma)$ be an $L^{2}$-normalized cusp form of weight $k$ and character $\chi$ for $\Gamma$. Then, we have the bound
(45) $\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\big{(}\|f(\tau)\|_{\mathrm{Pet}}^{2}\,\mathrm{Im}(\tau)^{\frac{1}{2}}\big{)}=O_{\Gamma_{0}}(k^{2}),$
where the implied constant depends on $\Gamma_{0}$.
@@
Since the width of the cusp $p_{1}=i\infty$ of $\Gamma$ equals $1$, the neighborhood $\mathcal{F}_{1}^{Y}$ of $p_{1}$ is contained in the fundamental domain $\mathcal{F}_{\Gamma_{0}}$, provided that $Y>1$. For a given $L^{2}$-normalized $f\in S_{k,\chi}(\Gamma)$, we then write
$\|f(\tau)\|_{\mathrm{Pet}}^{2}\,\mathrm{Im}(\tau)^{\frac{1}{2}}=\bigg{|}\frac{f(\tau)}{e^{2\pi i\tau}}\bigg{|}^{2}\cdot\frac{\eta^{k+1/2}}{e^{4\pi\eta}}.$
Now, the function $|f(\tau)/e^{2\pi i\tau}|^{2}$ is subharmonic and bounded in the neighborhood $\mathcal{F}_{1}^{Y}$ and thus takes its maximum on the boundary $\partial\mathcal{F}_{1}^{Y}$ by the maximum principle for subharmonic functions. Furthermore, by elementary calculus, we find that the function $\eta^{k+1/2}/e^{4\pi\eta}$ attains its maximum at
$\eta_{0}=\frac{k+1/2}{4\pi},$
and is strictly monotonically decreasing for $\eta>\eta_{0}$. By assuming for the moment that $k>12$ and by choosing
$Y\coloneqq\eta_{0}=\frac{k+1/2}{4\pi}>1,$
we derive from the arguments used in the proof of Theorem 3.5, in particular from the bound (LABEL:thm6-eqn4), that
$\sup_{\tau\in\mathcal{F}_{\Gamma_{0}}}\big{(}\|f(\tau)\|_{\mathrm{Pet}}^{2}\,\mathrm{Im}(\tau)^{\frac{1}{2}}\big{)}=\sup_{\tau\in\mathcal{F}_{1}^{\eta_{0}}}\big{(}\|f(\tau)\|_{\mathrm{Pet}}^{2}\,\eta^{\frac{1}{2}}\big{)}$
$\qquad\quad=\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\big{(}\|f(\tau)\|_{\mathrm{Pet}}^{2}\,\eta^{\frac{1}{2}}\big{)}\leq\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\big{(}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\,\eta^{\frac{1}{2}}\big{)}$
(46) $\qquad\quad\leq k^{\frac{3}{2}}\bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\sum_{\gamma\in\Gamma\setminus\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\bigg{)}.$
Finally, again using the arguments given in the proof of Theorem 3.5, shows that
(47) $\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\sum_{\gamma\in\Gamma\setminus\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O_{\Gamma_{0}}(1),$
(48) $\sup_{\tau\in\partial\mathcal{F}_{1}^{\eta_{0}}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O_{\Gamma_{0}}\big{(}k^{\frac{1}{2}}\big{)}.$
Combining the estimate (46) with the bounds (47) and (48), establishes the proof of the lemma for $k>12$. The finitely many remaining cases $k\in\{5,\ldots,12\}$ can now be obviously added to the statement by possibly enlarging the implied constant. ∎
Lemma 4.3.
For the index of the congruence subgroup
$\Gamma_{1}=\Gamma_{0,1}(4m)=\bigg{\{}\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{0}\,\bigg{|}\,c\equiv 0,\,d\equiv 1\mod 4m\bigg{\}}$
in $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, we have
$[\Gamma_{0}:\Gamma_{1}]=O_{\epsilon}\big{(}m^{2+\epsilon}\big{)},$
where the implied constant depends on the choice of $\epsilon>0$.
@@
The group $\Gamma_{0,1}(4m)$ fits into the short exact sequence
$0\longrightarrow\Gamma_{0,1}(4m)\longrightarrow\Gamma_{0}(4m)\longrightarrow(\mathbb{Z}/4m\mathbb{Z})^{\times}\longrightarrow 0,$
where the homomorphism from $\Gamma_{0}(4m)$ to $(\mathbb{Z}/4m\mathbb{Z})^{\times}$ is given by the assignment $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}\mapsto d$. From this we conclude that
$[\Gamma_{0}:\Gamma_{1}]$ $=[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]\cdot[\Gamma_{0}(4m):\Gamma_{0,1}(4m)]$
$=[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]\cdot\big{|}(\mathbb{Z}/4m\mathbb{Z})^{\times}\big{|}=[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]\cdot\varphi(m),$
where $\varphi(\cdot)$ denotes Euler’s $\varphi$-function. Since we trivially have $\varphi(m)\leq m$ and since it is well known that
$[\mathrm{SL}_{2}(\mathbb{Z}):\Gamma_{0}(4m)]=O_{\epsilon}\big{(}m^{1+\epsilon}\big{)}$
with an implied constant depending on the choice of $\epsilon>0$, the claim follows. ∎